### v1

@@ 5 Experiments
5.1 Experimental Setup
Given the narrow scope of few-shot skeleton-based action synthesis for HAR, we position ourselves in the same experimental setup proposed by [fukushi2024fewshot] to contextualize our results. We define our source pre-train dataset ($\mathcal{D}^{P}$) as HumanML3D [t2m_humanml] and the target ($\mathcal{D}^{T}$) as a subset of NTU RGB+D 120 [shahroudy2016nturgbdlargescale], analyzing actions: ’running on spot’ (A099), ’side kick’ (A102), and ’stretch on self’ (A104) from cross subject benchmark. Given the simplicity of these actions, generalist T2M models can reasonably generate them from broad text prompts. However, this setting precisely tests the challenge of capturing fine, class-discriminative kinematics critical for HAR. We choose these actions to evaluate KineMIC’s ability to close this kinematic specificity gap. Following the few-shot protocol, we randomly select only 10 samples per action class (30 total) as our support set ($\mathcal{D}^{\bar{T}}$) for framework training. Further details in supplementary material.
@@ 5.1 Experimental Setup
Figure 4: Qualitative Comparison of Generated Motions. The three columns depict the three target action classes. The first row shows corresponding real ground truth samples from NTU RGB+D 120 extracted through VIBE [vibe] (in orange), while following rows show diverse samples generated by our KineMIC model (in blue).
@@ 5.1.1 Dataset and preprocessing
Following [fukushi2024fewshot], 3D skeletons are re-estimated from RGB with VIBE [vibe]. To align NTU-derived target data with the HumanML3D pretraining [t2m_humanml] we apply two preprocessing steps: (1) drop hand joints from SMPL skeletons to match joint topology; (2) downsample sequences from 30→20 fps. All skeletons then follow the HumanML3D normalization pipeline: root centering, feet height set to zero, initial pose oriented to +Z, and joint-length normalization across frames. Motions are represented with the 263‑dim vector used in MDM [tevet2022humanmotiondiffusionmodel], concatenating root-relative joint positions, 6D joint rotations, joint velocities, and binary foot‑contact flags.
@@ 5.1.2 Implementation Details
The pre-trained model $G^{P}$ is an MDM 8-block transformer encoder [tevet2022humanmotiondiffusionmodel] trained on HumanML3D [t2m_humanml]. Our target model $G^{T}$ shares the identical architecture and weights of $G^{P}$, with the sole addition of an action embedding layer to its conditioning workflow (Fig. 2) and the MIC module. We fine-tune $G^{T}$ using LoRA [hu2021loralowrankadaptationlarge] (rank=16, $\alpha$=32, dropout=0.1) applied to all transformer layers, training only the low-rank matrices, the MIC module and the new action embedding layer. All other optimization objective $\lambda$ parameters are set to $1.0$. For the soft positive search, we set the number of nearest neighbors $k=250$ and a temperature of $\tau=0.07$. This $k$ value was chosen to effectively balance diversity and semantic relevance in the drawn samples. Search is pre-computed once, resulting in an effective training dataset size of $\textit{{(Number of Shots} + k)}\times\textit{{(Number of Classes)}}$. We employ classifier-free guidance with guidance scale of 2.5 [tevet2022humanmotiondiffusionmodel], dropping the Action and Text conditioning modalities each with independent probability of 0.1 during training. Models are trained for $5000$ steps using AdamW optimizer (LR $2\cdot 10^{-5}$) with $100$ diffusion steps and cosine noise scheduling. We apply gradient clipping (norm $1.0$) for early training stability. Each training step processes all $\mathcal{D}^{\bar{T}}$ samples (30 total), pairing each $(\mathbf{x}^{\bar{T}},y)$ with a randomly chosen soft-pair $(\mathbf{x}^{P},w)$ drawn from the top-$k$ matches in $\mathcal{D}^{P}$ relative to class $y$.
@@ 5.1.3 Evaluation Protocol
Our evaluation focuses on maximizing synthetic data utility for HAR, centered on recognition accuracy, Diversity (motion distribution variability) and MultiModality (per-prompt variance). We employ a ST-GCN classifier [yan2018spatialtemporalgraphconvolutional] for all downstream HAR and generative evaluations, following PYSKL practices and implementation [duan2022pysklgoodpracticesskeleton]. Each training set is composed of 30 real samples augmented with 1152 synthetically generated motions, both uniformly distributed per class. All experiments are repeated five times with different seeds. For comparison with prior work [fukushi2024fewshot], we report the median top-1 accuracy; internal analyses use mean and standard deviation. Further details in supplementary material.
@@ 5.2 Baselines and Prior Work Comparison
In Table 1 we compare our proposed KineMIC framework against key prior work and critical baselines to contextualize performance. First, we establish a foundational baseline using the pre-trained MDM model to generate motions directly from general text prompts describing our target classes (e.g., converting class ’side kick’ to ’a person does a side kick’), without any fine-tuning. This achieves 83.9%, confirming that the generalist MDM already possesses substantial motion knowledge relevant to our chosen HAR actions and providing a strong starting point. However, training MDM from random initialization (73.9%) or fine-tuning from a $\mathcal{D}^{P}$ checkpoint with LoRA (75.5%) both result in significantly lower performance. We attribute this to severe overfitting on the extremely limited few-shot data, causing the diffusion model to collapse. In stark contrast, our KineMIC framework better leverages pre-training knowledge while preventing this overfitting, achieving 86.2%. This gets close to the performance reported by [fukushi2024fewshot], but through a substantially different approach that leverages T2M diffusion priors rather than GAN-based regularization. Relative to real-data-only training on $\mathcal{D}^{\bar{T}}$ (63.1%), KineMIC delivers a substantial +23.1 percentage point improvement. Crucially, Figure 5 demonstrates scalability: real-data-only training shows poor performance and high variance in low-data regimes, while KineMIC augmentation consistently improves accuracy and asymptotically approaches full $\mathcal{D}^{T}$ training performance.
Table 1: Comparative Performance on Few-Shot HAR. Median top-1 accuracy on the NTU RGB+D 120 dataset, following the evaluation protocol defined by the prior work of Fukushi et al. [fukushi2024fewshot]. Results marked with $\dagger$} are reported directly from [fukushi2024fewshot]. The up-arrow ($\uparrow$) indicates that higher is better.
@@
Prior Work Real data only (30 samples) 58.4
ACTOR [petrovich2021actionconditioned3dhumanmotion] 73.6
Kinetic-GAN [degardin2021generativeadversarialgraphconvolutional] 81.7
Fukushi et al. [fukushi2024fewshot] 86.4
Our Analysis Real data only (30 samples) 63.1
@@ 5.4 Qualitative Evaluation
We conduct a comprehensive visual assessment of 75 generated samples from our best KineMIC model, alongside 75 outputs from the pre-trained MDM baseline and all 30 real $\mathcal{D}^{\bar{T}}$ samples (see supplementary material). Figure 4 presents select examples, demonstrating KineMIC’s ability to produce motions coherent with target kinematics while exhibiting rich, semantically meaningful variations, as supported by ablation gains in accuracy and diversity (Tab. 2). This diversity appears across action classes. For the ’side kick’ class, KineMIC generates motions ranging from $\mathcal{D}^{\bar{T}}$ replicas to dynamic combat-style kicks, demonstrating nuanced understanding beyond the sparse $\mathcal{D}^{\bar{T}}$ data. The ’stretch on self’ action produces enriched variations including stretches with hands behind the head or lateral torso tilts, while ’run on spot’ yields natural pacing and arm swing differences. Unlike pre-trained MDM’s generic variability, KineMIC outputs more consistently align with desirable $\mathcal{D}^{T}$ high level kinematics.
@@ 5.4.1 Motion Composition
KineMIC also exhibits an emergent compositional behavior when guided by a text prompt that differs from its action-conditioning class. When conditioned on ’stretch on self’ action id but prompted with the text ’a person is jumping’, for example, the model generates a plausible animation of a figure ’jumping while stretching its arms overhead’ (Fig. 6). We observe this property to be most successful when the two controls are not kinematically conflicting: combining an upper-body dominant action (’stretch on self’) with a lower-body one (’jumping’) produces coherent motions, whereas combining two lower-body dominant actions (’running on spot’ and ’jumping’) often leads to unstable results. This compositional ability likely emerges from the synergy between our dual-conditioning scheme (Fig. 2) and LoRA, which together help preserve both the fine-tuned action behavior and the general capabilities of the pre-trained model, in line with observations from [sawdayee2025dancelikechickenlowrank]. Although this behavior is not fully robust and a detailed analysis is beyond the scope of this work, it points to a promising avenue for future exploration.
@@ 6 Conclusions, limitations and future directions
In this work, the challenge of few-shot action synthesis for HAR using T2M priors is formally introduced and addressed. The study shows that T2M generative models can provide a strong baseline, yet often lack the kinematic specificity required for specialized, atomic action classes. To tackle these limitations, KineMIC, a teacher–student framework that adapts a pre-trained diffusion model using as few as 10 samples per class, is introduced. The main technical contributions are a soft positive search strategy that leverages a shared semantic space to retrieve relevant motion knowledge from the source domain, combined with a kinematic mining strategy. Empirically, KineMIC improves downstream classifier performance on the NTU RGB+D 120 few-shot benchmark [fukushi2024fewshot]. The key finding is that T2M priors represent a viable approach for few-shot HAR: naive fine-tuning leads to collapse (73.9-75.5%), whereas KineMIC’s context-aware mining delivers stable +23.1 percentage point gains over real-data-only baselines (63.1% → 86.2%).
Despite these advances, KineMIC has conceptual and practical limitations that point to future research directions. The primary constraint lies in the core assumption that semantic correspondence is a reliable proxy for kinematic relevance during the mining process. This assumption does not always hold; for instance, text encoders may assign high similarity to ‘punch’ and ‘kick’ due to shared concepts such as ‘fighting’, despite their substantial kinematic disparity, which increases the importance of effective filtering. As a result, robustness is intrinsically linked to (a) the scale and diversity of the T2M source data and (b) the specificity of the target actions, implying a practical limit when mining highly complex or novel atomic movements. Furthermore, prompt augmentation strategies were not explored. Integrating such techniques is expected to enhance the semantic richness of the target conditioning, a critical factor for fully exploiting the potential of T2M priors.

### v2

@@ 5 Experiments
Figure 4: Qualitative Comparison of Generated Motions. The three columns depict the three target action classes. The first row shows corresponding real ground truth samples from NTU RGB+D 120 extracted through VIBE [19] (in orange), while following rows show diverse samples generated by our KineMIC model (in blue).
5.1 Experimental Setup
Given the narrow scope of few-shot skeleton-based action synthesis for HAR, we position ourselves in the same experimental setup proposed by [12] to contextualize our results. We define our source pre-train dataset ($\mathcal{D}^{P}$) as HumanML3D [14] and the target ($\mathcal{D}^{T}$) as a subset of NTU RGB+D 120 [21], analyzing actions: ’running on spot’ (A099), ’side kick’ (A102), and ’stretch on self’ (A104) from cross subject benchmark. Given the simplicity of these actions, generalist T2M models can reasonably generate them from broad text prompts. However, this setting precisely tests the challenge of capturing fine, class-discriminative kinematics critical for HAR. We choose these actions to evaluate KineMIC’s ability to close this kinematic specificity gap. Following the few-shot protocol, we randomly select only 10 samples per action class (30 total) as our support set ($\mathcal{D}^{\bar{T}}$) for framework training. Further details in supplementary material.
@@ 5.1.1 Dataset and preprocessing
Following [12], 3D skeletons are re-estimated from RGB with VIBE [19]. To align NTU-derived target data with the HumanML3D pretraining [14] we apply two preprocessing steps: (1) drop hand joints from SMPL skeletons to match joint topology; (2) downsample sequences from 30→20 fps. All skeletons then follow the HumanML3D normalization pipeline: root centering, feet height set to zero, initial pose oriented to +Z, and joint-length normalization across frames. Motions are represented with the 263‑dim vector used in MDM [35], concatenating root-relative joint positions, 6D joint rotations, joint velocities, and binary foot‑contact flags.
@@ 5.1.2 Implementation Details
The pre-trained model $G^{P}$ is an MDM 8-block transformer encoder [35] trained on HumanML3D [14]. Our target model $G^{T}$ shares the identical architecture and weights of $G^{P}$, with the sole addition of an action embedding layer to its conditioning workflow (Fig. 2) and the MIC module. We fine-tune $G^{T}$ using LoRA [16] (rank=16, $\alpha$=32, dropout=0.1) applied to all transformer layers, training only the low-rank matrices, the MIC module and the new action embedding layer. All other optimization objective $\lambda$ parameters are set to $1.0$. For the soft positive search, we set the number of nearest neighbors $k=250$ and a temperature of $\tau=0.07$. This $k$ value was chosen to effectively balance diversity and semantic relevance in the drawn samples. Search is pre-computed once, resulting in an effective training dataset size of $\textit{{(Number of Shots} + k)}\times\textit{{(Number of Classes)}}$. We employ classifier-free guidance with guidance scale of 2.5 [35], dropping the Action and Text conditioning modalities each with independent probability of 0.1 during training. Models are trained for $5000$ steps using AdamW optimizer (LR $2\cdot 10^{-5}$) with $100$ diffusion steps and cosine noise scheduling. We apply gradient clipping (norm $1.0$) for early training stability. Each training step processes all $\mathcal{D}^{\bar{T}}$ samples (30 total), pairing each $(\mathbf{x}^{\bar{T}},y)$ with a randomly chosen soft-pair $(\mathbf{x}^{P},w)$ drawn from the top-$k$ matches in $\mathcal{D}^{P}$ relative to class $y$.
@@ 5.1.3 Evaluation Protocol
Our evaluation focuses on maximizing synthetic data utility for HAR, centered on recognition accuracy, Diversity (motion distribution variability) and MultiModality (per-prompt variance). We employ a ST-GCN classifier [38] for all downstream HAR and generative evaluations, following PYSKL practices and implementation [9]. Each training set is composed of 30 real samples augmented with 1152 synthetically generated motions, both uniformly distributed per class. Tests denoting "real-data-only" refer to experiments using only real samples ($\mathcal{D}^{\bar{T}}$) without synthetic augmentation. All experiments are repeated five times with different seeds. For comparison with prior work [12], we report the median top-1 accuracy; internal analyses use mean and standard deviation. Further details in supplementary material.
@@ 5.2 Baselines and Prior Work Comparison
In Table 1 we compare our proposed KineMIC framework against key prior work and critical baselines to contextualize performance. First, we establish a foundational baseline using the pre-trained MDM model to generate motions directly from general text prompts describing our target classes (e.g., converting class ’side kick’ to ’a person does a side kick’), without any fine-tuning. This achieves 83.9%, confirming that the generalist MDM already possesses substantial motion knowledge relevant to our chosen HAR actions and providing a strong starting point. However, training MDM from random initialization (73.9%) or fine-tuning from a $\mathcal{D}^{P}$ checkpoint with LoRA (75.5%) both result in significantly lower performance. We attribute this to severe overfitting on the extremely limited few-shot data, causing the diffusion model to collapse. In stark contrast, our KineMIC framework better leverages pre-training knowledge while preventing this overfitting, achieving 86.2%. This gets close to the performance reported by [12], but through a substantially different approach that leverages T2M diffusion priors rather than GAN-based regularization. Relative to real-data-only training on $\mathcal{D}^{\bar{T}}$ samples (63.1%), KineMIC delivers a substantial +23.1 percentage point improvement. Crucially, Figure 5 demonstrates scalability: real-data-only training shows poor performance and high variance in low-data regimes, while KineMIC augmentation consistently improves accuracy and asymptotically approaches full $\mathcal{D}^{T}$ training performance.
Table 1: Comparative Performance on Few-Shot HAR. Median top-1 accuracy on the NTU RGB+D 120 dataset, following the evaluation protocol defined by the prior work of Fukushi et al. [12]. Results marked with $\dagger$} are reported directly from [12]. The up-arrow ($\uparrow$) indicates that higher is better.
@@
Prior Work Real-data-only (30 samples) 58.4
ACTOR [25] 73.6
Kinetic-GAN [5] 81.7
Fukushi et al. [12] 86.4
Our Analysis Real-data-only (30 samples) 63.1
@@ 5.4 Qualitative Evaluation
We conduct a comprehensive visual assessment of 90 generated samples from our best KineMIC model, alongside 90 outputs from the pre-trained MDM baseline and all 30 real $\mathcal{D}^{\bar{T}}$ samples (see supplementary material). Figure 4 presents few examples, demonstrating KineMIC’s ability to produce motions coherent with target kinematics while exhibiting rich, semantically meaningful variations, as supported by ablation gains in accuracy and diversity (Tab. 2). For the ’side kick’ class, KineMIC generates motions ranging from $\mathcal{D}^{\bar{T}}$ replicas to dynamic combat-style kicks, demonstrating nuanced understanding beyond the sparse $\mathcal{D}^{\bar{T}}$ data. The ’stretch on self’ action produces enriched variations including stretches with hands behind the head or lateral torso tilts, while ’run on spot’ yields natural pacing and arm swing differences. Unlike pre-trained MDM’s generic variability, KineMIC outputs more consistently align with desirable $\mathcal{D}^{T}$ high level kinematics.
@@ 5.4.1 Motion Composition
KineMIC exhibits an emergent compositional behavior when guided by a text prompt that differs from its action-conditioning class. When conditioned on ’stretch on self’ action id but prompted with the text ’a person is jumping’, for example, the model generates a plausible animation of a figure ’jumping while stretching arms overhead’ (Fig. 6). We observe this property to be most successful when the two controls are not kinematically conflicting: combining an upper-body dominant action (’stretch on self’) with a lower-body one (’jumping’) produces coherent motions, whereas combining two lower-body dominant actions (’running on spot’ and ’jumping’) often leads to unstable results. This compositional ability likely emerges from the synergy between our dual-conditioning scheme (Fig. 2) and LoRA, which together help preserve both the fine-tuned action behavior and the general capabilities of the pre-trained model, in line with observations from [30]. Although this behavior is not fully robust and a detailed analysis is beyond the scope of this work, it points to a promising avenue for future exploration.
@@ 6 Conclusions, limitations and future directions
This work introduces and addresses the challenge of few-shot action synthesis for HAR using T2M priors. We show that while T2M models provide a strong baseline, they lack the kinematic specificity for specialized action classes. To resolve this, KineMIC adapts pre-trained diffusion models using only 10 samples per class via a soft positive search in a shared semantic space and a kinematic mining strategy. Empirically, KineMIC improves downstream classifier performance on the NTU-120 benchmark [12]. While naive fine-tuning leads to collapse (73.9–75.5%), KineMIC’s context-aware mining delivers stable +23.1 percentage point gains over real-data-only baselines (63.1% $\rightarrow$ 86.2%), proving T2M priors are viable for few-shot HAR.
Despite these advances, KineMIC has conceptual and practical limitations that point to future research directions. The primary constraint lies in the core assumption that semantic correspondence is a reliable proxy for kinematic relevance during the mining process. This assumption does not always hold; for instance, text encoders may assign high similarity to ‘punch’ and ‘kick’ due to shared concepts such as ‘fighting’, despite their substantial kinematic disparity, which increases the importance of effective filtering. As a result, robustness is intrinsically linked to (a) the scale and diversity of the T2M source data and (b) the specificity of the target actions, implying a practical limit when mining highly complex and specific movements. Furthermore, prompt augmentation strategies were not explored. Integrating such techniques is expected to enhance the semantic richness of the target conditioning, a critical factor for fully exploiting the potential of T2M priors.
Declarations.
Use of LLMs: AI tools were used solely for language refinement and technical editing. The core scientific contributions are the work of the authors, who hold full responsibility for the paper’s content. Acknowledgments: This work was conducted by the first author during an internship at CESI LINEACT under the second author’s supervision.